# Recalibrating binary probabilistic classifiers

### v1

@@ Abstract
Recalibration of binary probabilistic classifiers to a target prior probability is an important task in areas like credit risk management. We analyse methods for recalibration from a distribution shift perspective. Distribution shift assumptions linked to the area under the curve (AUC) of a probabilistic classifier are found to be useful for the design of meaningful recalibration methods. Two new methods called parametric covariate shift with posterior drift (CSPD) and ROC-based quasi moment matching (QMM) are proposed and tested together with some other methods in an example setting. The outcomes of the test suggest that the QMM methods discussed in the paper can provide appropriately conservative results in evaluations with concave functionals like for instance risk weights functions for credit risk.
@@ 1 Introduction
Recalibration of posterior probabilities learned on a training dataset to a target prior probability on a test dataset is not a well-defined problem because there is more than one way to transform the original posterior probabilities such that the target is matched. Therefore, we are going to study the impact of the recalibration method on the values of certain functionals of the posterior probabilities as an additional criterion to identify meaningful solutions. The risk weight functions for the calculation of minimum required capital under the “internal ratings based (IRB) approach” of the Basel credit risk framework (BCBS [2]) are a primary example of such functionals. The findings of this paper can inform the choice of the recalibration method in credit risk management and similar contexts.
This paper is organised as follows: Section 2 puts the paper into the context of related work. Section 3 describes the technical details of the setting of the paper and the recalibration problem. In addition, it introduces the evaluation of the solutions with a concave functional as a criterion for assessing their appropriateness. Section 4 presents a number of methods for recalibration. Parametric CSPD (Section 4.4) and ROC-based QMM (Section 4.5) seem to be new. With an example in Section 5, we illustrate the dependence of the solutions to the recalibration on assumptions of distribution shift and identify some less reliable recalibration methods. Section 6 proposes a way forward to prudent recalibration and concludes the paper. Appendix 0.A provides additional technical details needed for the implementation of some of the methods discussed in Section 4.
@@ 2 Related work
Quantification (or class distribution estimation, CDE) is another related problem. See Esuli et al. [7] for a recent survey. The goal of CDE in the binary case is to estimate the porportion of positive labels in a test dataset without any information on the labels. The primary common feature of the CDE and recalibration problems is the dependence of the solutions upon assumptions on the type of distribution shift between the training and test datasets.
Covariate shift with posterior drift (CSPD) as introduced by Scott [17] turns out to be a useful assumption on the type of distribution for tackling the recalibration problem. See Sections 4.4 and 5 below.
@@ 3.1 Recalibration
$\eta_{Q}(X)\ =\ T\bigl{(}\eta_{P}(X)\bigr{)}$ (1)
appears quite natural but actually is rather strong. Indeed, by Theorems 32.5 and 32.6 of Devroye et al. [6], (1) is equivalent to $\eta_{P}(X)$ being sufficient for $X$ with respect to $Y$ under $Q$, i.e. $Q[Y=1\,|\,X]\ =\ Q[Y=1\,|\,\eta_{P}(X)]$. This sufficiency property may be interpreted as ‘the information provided by $\eta_{P}(X)$ about $Y$ is as good as the information by the whole set of features $X$ under the target distribution $Q$’.
@@ 3.2 Non-uniqueness of recalibration
However, assuming a distribution shift between source $P$ and target $Q$ which results in a constant posterior probability $\eta_{Q}(x)=q$ appears too restrictive in most real world environments. In Section 5 below, we demonstrate that assuming preservation of classification performance as measured by AUC (area under the curve, see Appendix 0.A.1 below) between source $P$ and target $Q$ strikes a sensible note between too restrictive and too tolerant assumptions for the estimation of $E_{Q}\bigl{[}C(\eta_{Q}(X))\bigr{]}$ under the target features distribution.
@@ 4 Approaches to recalibration
Any solution as in (1) to the recalibration problem together with the marginal feature distribution $Q(X)$ completely determines the target distribution $Q(X,Y)$. Since $Q(X,Y)\neq P(X,Y)$ in case $q\neq p$, any given solution specifies some distribution shift between the source and target distributions. Hence, in order to better understand the consequences of selecting a particular solution transformation $T$, it is natural to explore the solution approaches through the lens of distribution shift. In the following, we assume $10<\eta_{P}(x)<$ 1 for all $x\in\mathcal{X}$.
@@ 4.1 Recalibration under assumption of stretched or compressed covariate shift
(3a) could be called capped scaling of the source posterior probabilities. Obviously, (3a) implies that (1) is satisfied with $T(z)=\min(t\,z,1)$.
Recall that source distribution $P(X,Y)$ and target distribution $Q(X,Y)$ are related through covariate shift (in the sense of this paper as defined by Storkey [19]) if it holds that $P[Y=y\,|\,X]=Q[Y=y\,|\,X]$ for all $y\in\mathcal{Y}$ with probability $1$ both under $P$ and $Q$. Hence if $t>1$ in (3a), one might call the implied distribution shift stretched covariate shift. In case $t<1$ the induced distribution shift could be considered compressed covariate shift. In case $q>E_{Q}[\eta_{P}(X)]$, (3a) and (3b) imply $t>1$ and $\eta_{Q}(x)=1$ for all $x$ with $t\,\eta_{P}(x)\geq 1$. This may have the consequence of an unjustified increase of $AUC_{\eta_{Q}(X)}$ compared to $AUC_{\eta_{P}(X)}$ for the area under the curve AUC defined as in Appendix 0.A.1 below.
@@ 4.2 Recalibration under assumption of label shift
Source distribution $P(X,Y)$ and target distribution $Q(X,Y)$ are related through label shift (Lipton et al. [11]), previously called prior probability shift in the literature (Storkey [19]), if $P[X\in M\,|\,Y=y]=Q[X\in M\,|\,Y=y]$ for all measurable subsets $M\subset\mathcal{X}$ and $y\in\mathcal{Y}$.
Under the assumption of label shift, the target feature distribution $Q(X)$ is represented as
@@
for all measurable sets $M\subset\mathcal{X}$. If a prior probability $q=Q[Y=1]$ is given, then according to the posterior correction formula (Eq. (2.4) of Saerens et al. [16]), $\eta_{Q}$ is determined through
@@
for all $x\in\mathcal{X}$ with probability $1$ under $Q$. In the credit risk community, the use of (5) for recalibration is popular (Section “Calibrating to PDs” of Bohn and Stein [3], Section 3.1 of Ptak-Chmielewska and Kopciuszewski [13]) because it avoids the problem of $\eta_{Q}(X)$ potentially taking the value $1$ for large $\eta_{P}(X)$ which may be encountered with capped scaling as in (3a). Note that (5) implies (1) with $T(z)=\frac{\frac{p}{q}\,z}{\frac{p}{q}\,z+\frac{1-p}{1-q}\,(1-z)}$ being strictly increasing.
Under the label shift assumption, the following observation is well-known. Define AUC (area under the curve) as in Appendix 0.A.1 below. AUC is a popular measure of performance of a classifier, i.e. AUC is considered an appropriate measure of the classifier’s ability to predict the true class label of an instance. See Chen et al. [5] for related comments.
@@
Proposition 1 is a consequence of (10a) in Appendix 0.A.1 below as well as (1) and (5). According to Proposition 1, recalibration under the assumption of label shift leaves the implied performance under the target distribution – sometimes called discriminatory power in the credit risk management literature (e.g. Bohn and Stein [3]) – unchanged when compared to the implied performance under the source distribution. In particular, Proposition 1 implies a necessary criterion for distribution shift to be label shift. In case $AUC_{S}\neq AUC_{S^{\ast}}$ source distribution $P$ and target distribution $Q$ cannot be related through label shift.
@@ 4.3 Recalibration under assumption of factorizable joint shift (FJS)
According to He at al. [9] and Tasche [23], the source distribution $P(X,Y)$ and the target distribution $Q(X,Y)$ are related through factorizable joint shift (FJS) if there are functions $g:\mathcal{X}\to[0,\infty)$ and $b:\mathcal{Y}\to[0,\infty)$ such that $(x,y)\mapsto g(x)\,b(y)$ for $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ is a density of $Q(X,Y)$ with respect to $P(X,Y)$, i.e. it holds that [^1] The indicator function $\mathbf{1}_{A}$ is defined by $\mathbf{1}_{A}(a)=1$ if $a\in A$ and $\mathbf{1}_{A}(a)=0$ if $a\notin A$. $Q[(X,Y)\in M]=E_{P}\bigl{[}\mathbf{1}_{M}(X,Y)\,g(X)\,b(Y)\bigr{]}$ for all measurable sets $M\subset\mathcal{X}\times\mathcal{Y}$.
By Corollary 4 of Tasche [23] and subject to mild technical conditions, under FJS the joint target distribution $Q(X,Y)$ is given by the target feature distribtuion $Q(X)$ and
@@
Note that recalibration of $\eta_{P}$ to a target prior $q$ under the assumption of FJS works for arbritrary target feature distributions $Q(X)$. This is in stark contrast to recalibration under the assumption of label shift which only works when assuming that $Q(X)$ is given by (4). However, recalibration under the label shift assumption entails $AUC_{\eta_{Q}(X)}=AUC_{\eta_{P}(X)}$. The example in Section 5 below shows that AUC preservation is not in general true for recalibration under an FJS assumption.
(6a) implies (1) with $T(z)=T_{\varrho}(z)=\frac{\frac{p}{q}\,z}{\frac{p}{q}\,z+\frac{1}{\varrho}\,\frac{1-p}{1-q}\,(1-z)}$ being strictly increasing. If $\varrho$ happens to take the value $1$, at first glance we are back in the context of label shift as in Section 4.2 above. However, as (4) need not hold true under the FJS assumption, in such cases there cannot be label shift. The type of shift modelled instead is called ‘invariant density ratio’ shift (Tasche [21]), defined as the special case of FJS with $\varrho=1$.
@@ 4.4 Recalibration under assumption of covariate shift with posterior drift (CSPD)
In Sections 4.1, 4.2 and 4.3, we identified the transformation $T$ of (1) after we had described a recalibration method designed under assumption of a certain type of distribution shift (slightly modified covariate shift, label shift, or FJS). In contrast, in this section, we define $T$ and use it to characterise the type of distribution shift implied by the recalibration method.
According to Scott [17], the source distribution $P(X,Y)$ and the target distribution $Q(X,Y)$ are related through covariate shift with posterior drift (CSPD) if there is a strictly increasing transformation $T$ such that (1) holds true.
As mentioned in Section 3.1, (1) implies that $\eta_{P}(X)$ is sufficient for $X$ with respect to $Y$ under the target distribution $Q$. Since under CSPD the transformation $T$ is strictly increasing, (1) also implies that $\eta_{Q}(X)$ and $\eta_{P}(X)$ are strongly comonotonic (Tasche [22]). As a consequence, Kendall’s $\tau$ and Spearman’s rank correlation both take the maximum value $1$ when applied to $(\eta_{Q}(X),\eta_{P}(X))$. This suggests that CSPD is a strong assumption which might be less often true than one would hope for. Nonetheless, comonotonicity of score and posterior probability is a common assumption in the literature. Chen et al. [5] call this assumption the rationality assumption.
If labels are available in the test dataset, under the CSPD assumption the transformation $T$ of (1) can be determined by means of isotonic regression. However, the general assumption for this paper is that there are no label observations for the instances in the test dataset. Therefore, we are going to apply a moment matching approach instead (quasi moment matching, QMM).
To be able to do so, we consider parametric CSPD where the transformation $T$ of (1) is specified as follows through some strictly increasing and continuous distribution function $F$ on the real line and parameters $a,b\in\mathbb{R}$:
@@
  • Standard normal distribution (inverse probit) function: $F(x)=\Phi(x)$, $x\in\mathbb{R}$. We refer to this choice of $F$ as normal CSPD.
@@
with $T_{a,b}$ as in (7a) or (7b) and AUC defined in Appendix 0.A.1. More precisely, in (7c), $AUC_{\eta_{P}(X)}$ is computed with respect to $P(X)$ while $AUC_{T_{a,b}(\eta_{P}(X))}$ is computed as implied AUC with respect to $Q(X)$. See also Appendix 0.A.1 for the different ways to compute AUC.

### v2

@@ Abstract
Recalibration of binary probabilistic classifiers to a target prior probability is an important task in areas like credit risk management. We analyse methods for recalibration from a distribution shift perspective. Distribution shift assumptions linked to the area under the curve (AUC) of a probabilistic classifier are found to be useful for the design of meaningful recalibration methods. Two new methods called parametric covariate shift with posterior drift (CSPD) and ROC-based quasi moment matching (QMM) are proposed and tested together with some other methods in an example setting. The outcomes of the test suggest that the QMM methods discussed in the paper can provide appropriately conservative results in evaluations with concave functions like for instance risk weights functions for credit risk.
@@ 1 Introduction
This is a common situation in credit risk management, see for instance Chapter 4 of Bohn and Stein [3]. So called probabilities of default (PDs) are estimated on a training sample with observations of solvent and defaulted borrowers and must be recalibrated before being evaluated for the borrowers in a live portfolio. The future solvency states of these borrowers in general are unknown but typically an estimate of the proportion of borrowers who are going to default is available. Sometimes, such estimates are conservative, i.e. they are likely to significantly overestimate the proportion of defaulters. Conservatism of the estimates can be a regulatory requirement or it could be part of stress testing exercises intended to assess the impact of unfavourable economic circumstances on the portfolio.
Recalibration of posterior probabilities learned on a training dataset to a target prior probability on a test dataset is not a well-defined problem because there is more than one way to transform the original posterior probabilities such that the target is matched. Therefore, we are going to study the impact of the recalibration method on the values of concave or nearly concave functions of the posterior probabilities as an additional criterion to identify meaningful solutions. The risk weight functions for the calculation of minimum required capital under the “internal ratings based (IRB) approach” of the Basel credit risk framework (BCBS [2]) are a primary example of such functions. The findings of this paper can inform the choice of the recalibration method in credit risk management and similar contexts.
This paper is organised as follows: Section 2 puts the paper into the context of related work. Section 3 describes the technical details of the setting of the paper and the recalibration problem. In addition, it introduces the evaluation of the solutions with a concave function as a criterion for assessing their appropriateness. Section 4 presents a number of methods for recalibration. Parametric CSPD (Section 4.4) and ROC-based QMM (Section 4.5) seem to be new. With an example in Section 5, we illustrate the dependence of the solutions to the recalibration problem on assumptions of distribution shift and identify some less reliable recalibration methods. Section 6 proposes a way forward to prudent recalibration and concludes the paper. Appendix 0.A provides additional technical details needed for the implementation of some of the methods discussed in Section 4.
@@ 2 Related work
Quantification (or class distribution estimation, CDE) is another related problem. See Esuli et al. [8] for a recent survey. The goal of CDE in the binary case is to estimate the proportion of positive labels in a test dataset without any information on the labels. The primary common feature of the CDE and recalibration problems is the dependence of the solutions upon assumptions on the type of distribution shift between the training and test datasets. However, Remark 1 in Section 4.4 shows that the one-parameter version of the recalibration method ‘parametric CSPD’ may also be used for CDE.
Covariate shift with posterior drift (CSPD) as introduced by Scott [19] turns out to be a useful assumption on the type of distribution shift for tackling the recalibration problem. See Sections 4.4 and 5 below.
@@ 3.1 Recalibration
$\eta_{Q}(X)\ =\ T\bigl{(}\eta_{P}(X)\bigr{)}$ (1b)
appears quite natural but actually is rather strong. Indeed, by Theorems 32.5 and 32.6 of Devroye et al. [7], (1b) is equivalent to $\eta_{P}(X)$ being sufficient for $X$ with respect to $Y$ under $Q$, i.e. $Q[Y=1\,|\,X]\ =\ Q[Y=1\,|\,\eta_{P}(X)]$. This sufficiency property may be interpreted as ‘the information provided by $\eta_{P}(X)$ about $Y$ is as good as the information by the whole set of features $X$ under the target distribution $Q$’.
@@ 3.2 Non-uniqueness of recalibration
However, assuming a distribution shift between source $P$ and target $Q$ which results in a constant posterior probability $\eta_{Q}(x)=q$ appears too restrictive in most real world environments. In Section 5 below, we demonstrate that assuming preservation of classification performance as measured by AUC (Area Under the Curve, see next section) between source $P$ and target $Q$ strikes a sensible note between too restrictive and too tolerant assumptions for the estimation of $E_{Q}\bigl{[}C(\eta_{Q}(X))\bigr{]}$ under the target features distribution.
3.3 Area Under the Curve (AUC)
AUC (Area Under the Curve111‘Curve’ refers to ROC (Receiver Operating Characteristic). See Fawcett [9] and the references therein for the most common definitions of ROC and AUC.) is a popular measure of performance of a binary classifier, i.e. AUC is considered an appropriate measure of the classifier’s ability to predict the true class label of an instance. See Chen et al. [5] for related comments. Since AUC plays an important role in some of the recalibration methods discussed in the following sections, we present here the population-level (in contrast to sample-based) representations of AUC that are used in this paper.
$AUC_{S}\ =\ P^{\ast}[S_{1}>S_{0}]+\frac{1}{2}\,P^{\ast}[S_{1}=S_{0}],$ (3a)
However, $AUC_{S}$ can also be determined if the distribution $P(S)$ of $S$ and the posterior probabilities $\eta_{P}(S)=P[Y=1\,|\,S]$ are known. Then with $p=E_{P}[\eta_{P}(S)]$ it holds that [^2] The indicator function $\mathbf{1}_{A}$ is defined by $\mathbf{1}_{A}(a)=1$ if $a\in A$ and $\mathbf{1}_{A}(a)=0$ if $a\notin A$.
$\begin{split}P_{1}[S\in M]&\ =\ \frac{E_{P}\bigl{[}\eta_{P}(S)\,\mathbf{1}_{M}(S)\bigr{]}}{p}\qquad\text{and}\\ P_{0}[S\in M]&\ =\ \frac{E_{P}\bigl{[}(1-\eta_{P}(S))\,\mathbf{1}_{M}(S)\bigr{]}}{1-p}\end{split}$ (3b)
for all measurable sets $M\subset\boldsymbol{\mathcal{S}}$. Plug $P_{0}$ and $P_{1}$ from (3b) in (3a) to compute $AUC_{S}$.
To indicate the way $AUC_{S}$ is computed, in this paper we refer simply to AUC if $AUC_{S}$ is assumed to be computed or inferred by means of (3a) irrespectively of the origin of $P_{0}$ and $P_{1}$. We refer to implied AUC if $AUC_{S}$ is assumed to be computed by means of (3b) in combination with (3a). See Appendix 0.A.1 for a more detailed formula for implied AUC in the case of a discrete-valued score $S$.
@@ 4 Approaches to recalibration
Any solution as in (1b) to the recalibration problem together with the marginal feature distribution $Q(X)$ completely determines the target distribution $Q(X,Y)$. Since $Q(X,Y)\neq P(X,Y)$ in case $q\neq p$, any given solution specifies some distribution shift between the source and target distributions. Hence, in order to better understand the consequences of selecting a particular solution transformation $T$, it is natural to explore the solution approaches through the lens of distribution shift. In the following, we assume $10<\eta_{P}(x)<$ 1 for all $x\in\mathcal{X}$.
@@ 4.1 Recalibration under assumption of stretched or compressed covariate shift
(4a) could be called capped scaling of the source posterior probabilities. Obviously, (4a) implies that (1b) is satisfied with $T(\eta)=\min(t\,\eta,1)$.
Recall that source distribution $P(X,Y)$ and target distribution $Q(X,Y)$ are related through covariate shift (in the sense of Storkey [21]) if it holds that $P[Y=y\,|\,X]=Q[Y=y\,|\,X]$ for all $y\in\mathcal{Y}$ with probability $1$ both under $P$ and $Q$. Hence if $t>1$ in (4a), one might call the implied distribution shift stretched covariate shift.
In case $t<1$ the induced distribution shift could be considered compressed covariate shift. In case $q>E_{Q}[\eta_{P}(X)]$, (4a) and (4b) imply $t>1$ and $\eta_{Q}(x)=1$ for all $x$ with $t\,\eta_{P}(x)\geq 1$. This may have the consequence of an unjustified increase of $AUC_{\eta_{Q}(X)}$ compared to $AUC_{\eta_{P}(X)}$ for the area under the curve AUC defined as in Section 3.3 below.
@@ 4.2 Recalibration under assumption of label shift
Source distribution $P(X,Y)$ and target distribution $Q(X,Y)$ are related through label shift (Lipton et al. [12]), previously called prior probability shift in the literature (Storkey [21]), if $P[X\in M\,|\,Y=y]=Q[X\in M\,|\,Y=y]$ for all measurable sets $M\subset\mathcal{X}$ and $y\in\mathcal{Y}$.
Under the assumption of label shift, the target feature distribution $Q(X)$ can be represented as
@@
for all measurable sets $M\subset\mathcal{X}$. If a prior probability $q=Q[Y=1]$ is given, then according to the posterior correction formula (Eq. (2.4) of Saerens et al. [18]), $\eta_{Q}$ is determined through (recall $p=P[Y=1]$)
@@
for all $x\in\mathcal{X}$ with probability $1$ under $Q$. In the credit risk community, the use of (6) for recalibration is popular (Section “Calibrating to PDs” of Bohn and Stein [3], Section 3.1 of Ptak-Chmielewska and Kopciuszewski [15]) because it avoids the problem of $\eta_{Q}(X)$ potentially taking the value $1$ for large $\eta_{P}(X)$ which may be encountered with capped scaling as in (4a). Cramer [6] (Sections 6.2 and 6.3) pointed out that in the context of logistic regression the related special case of (6) was known at least since 1979. Note that (6) implies (1b) with $T(\eta)=\frac{\frac{p}{q}\,\eta}{\frac{p}{q}\,\eta+\frac{1-p}{1-q}\,(1-\eta)}$ strictly increasing in $\eta$.
Under the label shift assumption, the following observation is well-known. Define AUC (area under the curve) as in Section 3.3.
@@
Proposition 1 is a consequence of (3a) in Section 3.3 as well as (1b) and (6). According to Proposition 1, recalibration under the assumption of label shift leaves the implied performance under the target distribution – sometimes called discriminatory power in the credit risk management literature (e.g. Bohn and Stein [3]) – unchanged when compared to the implied performance under the source distribution. In particular, Proposition 1 implies a necessary criterion for distribution shift to be label shift. Accordingly, in case $AUC_{S}\neq AUC_{S^{\ast}}$ source distribution $P$ and target distribution $Q$ cannot be related through label shift.
@@ 4.3 Recalibration under assumption of factorizable joint shift (FJS)
According to He at al. [10] and Tasche [25], the source distribution $P(X,Y)$ and the target distribution $Q(X,Y)$ are related through factorizable joint shift (FJS) if there are functions $g:\mathcal{X}\to[0,\infty)$ and $b:\mathcal{Y}\to[0,\infty)$ such that $(x,y)\mapsto g(x)\,b(y)$ for $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ is a density of $Q(X,Y)$ with respect to $P(X,Y)$, i.e. it holds that $Q[(X,Y)\in M]=E_{P}\bigl{[}\mathbf{1}_{M}(X,Y)\,g(X)\,b(Y)\bigr{]}$ for all measurable sets $M\subset\mathcal{X}\times\mathcal{Y}$. See footnote 2 for the definition of the indicator $\mathbf{1}_{M}$.
By Corollary 4 of Tasche [25] and subject to mild technical conditions, under FJS the joint target distribution $Q(X,Y)$ is given by the target feature distribtuion $Q(X)$ and the class 1 posterior probability
@@
Note that recalibration of $\eta_{P}(X)$ to a target prior probability $q$ under the assumption of FJS works for arbritrary target feature distributions $Q(X)$. This is in stark contrast to recalibration under the assumption of label shift which only works when assuming that $Q(X)$ is given by (5). However, by Proposition 1 recalibration under the label shift assumption entails $AUC_{\eta_{Q}(X)}=AUC_{\eta_{P}(X)}$. The example in Section 5 below shows that AUC preservation is not in general true for recalibration under the FJS assumption.
(7a) implies (1b) with $T(\eta)=T_{\varrho}(\eta)=\frac{\frac{p}{q}\,\eta}{\frac{p}{q}\,\eta+\frac{1}{\varrho}\,\frac{1-p}{1-q}\,(1-\eta)}$ strictly increasing in $\eta$. If $\varrho$ happens to take the value $1$, at first glance we are back in the context of label shift as in Section 4.2 above. However, as (5) need not hold true under the FJS assumption, in such cases there cannot be label shift. The type of shift modelled instead is called ‘invariant density ratio’ shift (Tasche [23]), defined as the special case of FJS with $\varrho=1$.
@@ 4.4 Recalibration under assumption of covariate shift with posterior drift (CSPD)
In Sections 4.1, 4.2 and 4.3, we identified the transformation $T$ of (1b) after we had described a recalibration method designed under assumption of one of three types of distribution shift, namely slightly modified covariate shift, label shift, and FJS. In contrast, in this section, we define $T$ and use it to characterise the type of distribution shift implied by the recalibration method.
According to Scott [19], the source distribution $P(X,Y)$ and the target distribution $Q(X,Y)$ are related through covariate shift with posterior drift (CSPD) if there is a strictly increasing transformation $T$ such that (1b) holds true.
As mentioned in Section 3.1, (1b) implies that $\eta_{P}(X)$ is sufficient for $X$ with respect to $Y$ under the target distribution $Q$. Since under CSPD the transformation $T$ is strictly increasing, (1b) also implies that $\eta_{Q}(X)$ and $\eta_{P}(X)$ are strongly comonotonic (Tasche [24]). As a consequence, Kendall’s $\tau$ and Spearman’s rank correlation both take the maximum value $1$ when applied to $(\eta_{Q}(X),\eta_{P}(X))$. This suggests that CSPD is a strong assumption which might be less often true than one would hope for. Nonetheless, comonotonicity of score and posterior probability is a common assumption in the literature. Chen et al. [5] call this assumption the rationality assumption.
If labels are available in the test dataset, under the CSPD assumption the transformation $T$ of (1b) can be approximately determined by means of isotonic regression. However, the general assumption for this paper is that there are no label observations for the instances in the test dataset. Therefore, we are going to apply a moment matching approach instead (quasi moment matching, QMM).
To be able to do so, we consider parametric CSPD where the transformation $T$ of (1b) is specified as follows through some strictly increasing and continuous distribution function $F$ on the real line and parameters $a,b\in\mathbb{R}$:
@@
  • Standard normal distribution function (inverse probit): $F(x)=\Phi(x)$, $x\in\mathbb{R}$. We refer to this choice of $F$ as normal CSPD.
@@
with $T_{a,b}$ as in (8a) or (8b) and AUC defined in Section 3.3. More precisely, in (8c), $AUC_{\eta_{P}(X)}$ is computed with respect to $P(X,Y)$ while $AUC_{T_{a,b}(\eta_{P}(X))}$ is computed as implied AUC with respect to $Q(X)$. See Section 3.3 for the different ways to compute AUC.
@@
Remark 1
One-parameter CSPD as in (8a) with $b=0$ can be used for class distribution estimation (also called quantification), i.e. to determine an unknown class prior probability $q=Q[Y=1]$ under the target distribution. In this case, instead of solving (8c) for two parameters $a$ and $b$, the following equation is solved for parameter $a$ only:
$AUC_{\eta_{P}(X)}\ =\ AUC_{T_{a,0}(\eta_{P}(X))}.$ (9a)
Then the mean of the resulting posterior probability $\eta_{Q}(X)=T_{a,0}(\eta_{P}(X))$ under the target distribution feature distribution $Q(X)$ is computed to obtain an estimate $\widehat{q}$ of $q$:
$\widehat{q}\ =\ E_{Q}\bigl{[}T_{a,0}(\eta_{P}(X))\bigr{]}.$ (9b)
One-parameter CSPD may be interpreted as a modification of quantification under the assumption of covariate shift. In contrast to assuming covariate shift when AUC can differ between source distribution and target distribution, with one-parameter CSPD by construction source AUC and target AUC are equal.